You randomly select and weigh 26 samples of an allergy medication. The sample standard deviation is 1.4 milligrams. Assuming the weights are normally distributed, construct 95% confidence intervals for the population variance and standard deviation.
Answer #1

Solution :

Given that,

c = 0.95

s = 1.4

n = 26

At 95% confidence level the \chi ^{2} is ,

\alpha = 1 - 95% = 1 - 0.95= 0.05

\alpha / 2 = 0.05 / 2 = 0.01

\chi ^{2}\alpha /2,df = \chi ^{2} 0.025,25 = 40.65

and

\chi ^{2}1-\alpha /2,df = \chi ^{2} 0.975,18 = 13.12

Point estimate = s2 = 1.96

\chi2L = \chi 2\alpha/2,df = 40.65

\chi2R = \chi 21 - \alpha /2,df = 13.12

The 98% confidence interval for \sigma 2 is,

(n - 1)s2 / \chi 2\alpha/2 < \sigma 2 < (n - 1)s2 / \chi 21 - \alpha /2

( 25 *1.96) / 40.65< \sigma 2 < ( 25* 1.96) / 13.12

1.21< \sigma 2 < 3.73

(1.21, 3.73)

The 98% confidence interval for \sigma is,

s \sqrt{} (n-1) / \chi ^{2} \alpha /2,df < \sigma < s \sqrt{} (n-1) / \chi ^{2} 1-\alpha /2,df

1.4\sqrt{}( 26 - 1 ) / 40.65 < \sigma < 1.4\sqrt{}( 26- 1 ) / 13.12

1.10< \sigma < 1.93

( 1.10 , 1.93)